# qlty: handling large tensors in scientific imaging

### v1

@@ 1 Introduction
In computer vision applications within the experimental sciences, such as in the analysis of X-ray or Electron Tomography, Focused Ion Beam Scanning Electron Microscopy (FIB-SEM) or Adaptive Optics Lattice Light Sheet Microscopy (AO-LLSM) [1, 2, 3, 4] the use of deep learning networks for image processing and vision-related tasks has become increasingly important. These workflows often require substantial computational resources, as data set sizes easily contain over $10^{9}$ voxels. A significant challenge arises when training these networks, particularly with volumetric data, as the associated memory requirements for machine learning approaches on such datasets rapidly outrun the available capacity conventional hardware has to offer. For inference tasks, these issues arise as well, pushing the need for efficient data handling and processing techniques.
While distributed computing approaches, where data is shared across multiple GPUs, offer a potential solution, they are not always feasible. In scenarios involving consumer hardware or edge computing, where available resources may be significantly constrained compared to dedicated compute clusters, the problems are even more pressing. Moreover, in many scientific applications, particularly in segmentation tasks, the available labeled data can be sparse within large volumetric datasets. This sparsity presents an additional opportunity: regions of the data that lack annotations and lie outside the effective receptive field of the neural network contribute little to model training. Processing these unlabeled regions wastes valuable computational resources without improving model performance. Consequently, there is a pressing need for efficient tensor management techniques that can handle spatial data effectively in a resource-limited environment while also intelligently focusing computation on the most informative regions of sparsely annotated datasets.
@@
Out-of-core processing methods have emerged as a crucial approach to handling datasets that exceed available memory, offering a practical solution to the challenges posed by large volumetric data in scientific computer vision [5]. These techniques typically involve data chunking, where the dataset is divided into manageable portions that can be loaded and processed sequentially, allowing for the analysis of data that far exceeds the capacity of RAM or GPU memory. While effective in many data processing scenarios, existing out-of-core methods often lack the specific functionality required for complex computer vision tasks, particularly those involving deep learning models. A key requirement for a tool to aid in these scenarios is the ability to handle overlapping subtensors, both during the chunking process and when stitching processed results back together. This capability is crucial for maintaining spatial continuity and addressing edge effects in convolutional neural networks. The absence of a light-weight tool that could efficiently manage these tasks, while also integrating seamlessly with PyTorch’s tensor operations for computer vision applications, drove the development of new methodologies to fill this gap. Such a tool needs to be capable of handling large-scale datasets, providing robust solutions for tensor patching, manipulation, and stitching.
@@
The qlty toolkit addresses these challenges through a suite of specialized functions designed for tensor management in computer vision applications. By providing robust methods for tensor patching, manipulation, and stitching, qlty enables researchers and developers to work effectively with large-scale spatial data, even in resource-constrained environments. In this paper, we present the functionality of qlty and demonstrate its effectiveness through a segmentation on experimental data, illustrating how qlty can enable the processing of larger datasets than would otherwise be possible on limited hardware.
@@ 2 Purpose and Functionality
Figure 1: High level overview of qlty functionality. A - For generating training data with a smaller spatial footprint, qlty subsamples a sparsely annotated tensor $\mathcal{T}$ into smaller chunks $\mathcal{S}$ while discarding parts of $\mathcal{T}$ for which no training data is available. B - In inference scenarios, qlty iterates over data tensor $\mathcal{T}$ yielding manageable chunks of data that can be subjected to a neural network of choice. The resulting output is iterative placed back into a tensor of the right size.
A more detailed description of the functionality is given below.
@@ 2.1 Subsampling
Assume we have a PyTorch tensor $\mathcal{T}$ with shape $(N,C,Y,X)$. This data structure contains $N$ items of $C$ channel data of a two-dimensional spatial map with lengths $Y$ and $X$. When subsampling $\mathcal{T}$, the aim is to obtain tensor $\mathcal{S}$ of size $(M,C,Y_{w},X_{w})$, where $(Y_{w},X_{w})$ is referred to as the window size. The channel count $C$ remains untouched, but the resulting batch size $M$ will typically be not equal to $N$. Within qlty , the subsampling is done in a deterministic fashion by translating a window with a predefined step size across the input tensor, while deploying an step-back correction at the edges of tensor $\mathcal{T}$ to ensure that the last subsampled window has edges aligned with the input tensor. This process is outlined for a multi-channel 2D tensors in Algorithm 1 and extends analogously to 3D scenarios.
@@ 3 Example Usage
For inference, a larger window size is technically feasible because of lower memory requirements. Using a window size of 128, a step size of 108 and a border width of 10 were used. By setting the border weight to zero, the resulting stitching operation results in non-overlapping tensor of size $(108,108,108)$. In this manner, 231 subtensors were generated and run through an inference pass. Further qlty setting are found in Table 2.
@@ 4 Impact
The seamless integration of qlty with the PyTorch framework, streamlines machine learning workflows for scientific imaging. qlty’s subsampling and stitching functions can be easily incorporated within existing data analytics pipelines [9], circumventing the need to write ad-hoc functionality that addresses these issues. qlty thus reduces a barrier to utilize advanced image processing techniques in the broader scientific community.
qlty is available a pip installable package, or via https://github.com/phzwart/qlty.
@@ 5 Acknowledgements
We gratefully acknowledge the support of this work by the Laboratory Directed Research and Development Program of Lawrence Berkeley National Laboratory under US Department of Energy contract No. DE-AC02-05CH11231. Further support originated from the Center for Advanced Mathematics in Energy Research Applications funded via the Advanced Scientific Computing Research and the Basic Energy Sciences programs, which are supported by the Office of Science of the US Department of Energy (DOE) under contract No. DE-AC02-05CH11231, and from the National Institute of General Medical Sciences of the National Institutes of Health (NIH) under award 5R21GM129649-02. We acknowledge the use of resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility operated under Contract No. DE-AC02-05CH11231, under NERSC allocation m4055.

### v2

@@ Abstract
Current code version 0.2.0
Permanent link to code/repository used for this code version https://github.com/phzwart/qlty/releases/tag/v0.2.0
Permanent link to reproducible capsule https://codeocean.com/capsule/c5aabf63-6d0b-4e4b-abbc-99abe85870c8/
Legal code license BSD License
Code versioning system used git
Software code languages, tools and services used Python, Jupyter Notebook
Compilation requirements, operating environments and dependencies Python, pytorch, dask, einops
If available, link to developer documentation/manual https://qlty.readthedocs.io/
Support email for questions PHZwart@lbl.gov
Table 1: Code metadata
@@ 1 Introduction
In computer vision applications within the experimental sciences, such as in the analysis of X-ray or Electron Tomography, Focused Ion Beam Scanning Electron Microscopy (FIB-SEM), or Adaptive Optics Lattice Light Sheet Microscopy (AO-LLSM) [1, 2, 3, 4, 5], the use of deep learning networks for image processing and vision-related tasks has become increasingly important. A prime example of the use of deep learning techniques is in semantic segmentation [6, 7, 8, 9, 10], where pixel-level classifications in images are made so specific objects can be extracted, or properties of a sample can be derived [11, 12].
When these large datasets are inspected, visualized, or analyzed, they are done as a complete object. Manual annotation often results in labeling only part of an image, the full volume is still needed to contextualize the information during the annotation process as smaller cropped sections can make proper labeling difficult. Similarly, when extracting downstream conclusions after segmentation, the spatial relationships within the full object need to remain intact. While libraries for handling large [13] and very large [14] data volumes are available, these methods are not directly geared towards enabling deep learning, necessitating the development of specialized tools and techniques for handling, processing, and segmenting large datasets within deep learning workflows. During the training and inference of a deep learning workflow, the associated memory requirements for large datasets rapidly outruns the available capacity conventional hardware has to offer, pushing the need for efficient data handling and processing techniques.
@@
Out-of-core processing methods have emerged as a crucial approach to handling datasets that exceed available memory, offering a practical solution to the challenges posed by large volumetric data in scientific computer vision [14]. These techniques typically involve data chunking, where the dataset is divided into manageable portions that can be loaded and processed sequentially, allowing for the analysis of data that far exceeds the capacity of RAM or GPU memory. While effective in many data processing scenarios, existing out-of-core methods often lack the specific functionality required for complex computer vision tasks, particularly those involving deep learning models. A key requirement for a tool to aid in these scenarios is the ability to handle overlapping subtensors, both during the chunking process and when stitching processed results back together. This capability is crucial for maintaining spatial continuity and addressing edge effects in convolutional neural networks. The absence of a light-weight tool that can efficiently manage patching, manipulation and stitching, while also integrating seamlessly with PyTorch’s tensor operations for computer vision applications, drove the development to fill this gap.
@@
Here we present qlty, a lightweight python toolkit that addresses these challenges through a suite of specialized functions. By providing robust methods for tensor patching, manipulation, and stitching, qlty enables researchers and developers to work effectively with large-scale spatial data, even in resource-constrained environments. We will illustrate qlty’s functionality and demonstrate its effectiveness through a segmentation of experimental data, illustrating how qlty can enable the processing of larger datasets that would otherwise be challenging on limited hardware.
@@ 2 Purpose and Functionality
Figure 1: High level overview of qlty’s functionality, with specific methods depicted in red. A - For generating training data with a smaller spatial footprint, qlty subsamples a sparsely annotated tensor $\mathcal{T}$ into smaller chunks $\mathcal{S}$ while discarding parts of $\mathcal{T}$ for which no training data is available. B - In inference scenarios, qlty iterates over data tensor $\mathcal{T}$ yielding manageable chunks of data that can be subjected to a neural network of choice. The resulting output is iteratively placed back into a tensor of the right size.
@@ 2.1 Subsampling
Assume we have a PyTorch tensor $\mathcal{T}$ with shape $(N,C,Y,X)$. This data structure contains $N$ items of $C$ channel data of a two-dimensional spatial map with lengths $Y$ and $X$. When subsampling $\mathcal{T}$, the aim is to obtain tensor $\mathcal{S}$ of size $(M,C,Y_{w},X_{w})$, where $(Y_{w},X_{w})$ is referred to as the window size. The channel count $C$ remains untouched, but the resulting batch size $M$ will typically be not equal to $N$. Within qlty, subsampling is done in a deterministic fashion by translating a window with a predefined step size across the input tensor, while deploying an step-back correction at the edges of tensor $\mathcal{T}$ to ensure that the last subsampled window has edges aligned with the input tensor. This process is outlined for a multi-channel 2D tensor in Algorithm 1 and extends analogously to 3D scenarios. At present, 2D + time or 3D + time scenarios are not supported, but will be added in the future.
@@ 3 Example Usage
These trained networks were used to run inference across the full data volume. For inference, a larger window size is technically feasible because of lower memory requirements. Using a window size of 128, a step size of 108 and a border width of 10 were used. By setting the border weight to zero, the resulting stitching operation results in patching together non-overlapping tensor of size $(108,108,108)$. In this manner, 231 subtensors were generated and run through an inference pass. Further qlty setting are found in Table 3.
@@ 4 Impact
The practical impact of qlty as observed in the aforementioned applications [18, 19, 20], lies in its integration within standard workflows, allowing users or automated processes to adjust window and step sizes to easily manage GPU memory limitations. It is commonly embedded as a preprocessing step in segmentation or denoising pipelines, simplifying the process of running these workflows across different hardware architectures, such as RTX 3090s or A100 GPUs. While memory constraints are often addressed by adjusting batch sizes, breaking down tensors into smaller subtensors to fit within these limitations is typically more complex — qlty effectively streamlines this process.
The seamless integration of qlty within the PyTorch framework, streamlines machine learning workflows for scientific imaging. qlty’s subsampling and stitching functions can be easily incorporated within existing data analytics pipelines [18], circumventing the need to write ad-hoc functionality that addresses these issues. qlty thus reduces a barrier to utilize advanced image processing techniques in the broader scientific community.
5 Future Directions
By working closely with experimental scientists, a number of future development directions have been identified.
Supporting Fragmented Workflows: In many applications, the locations where data is generated, computation is performed, and where results are visualized are often different and distributed across various network locations. A promising future direction for qlty is to integrate it with technologies that allow seamless data access across these disparate networked environments. By combining qlty with distributed computing frameworks and remote data access protocols, we can facilitate efficient data handling and computation even in workflows fragmented across many locations, ensuring smooth operations regardless of where the data resides.
Supporting Image Pyramids: Image pyramids are powerful data structures for managing large datasets across different scales. Neural networks that leverage these pyramids can effectively decouple receptive field requirements across scale spaces, potentially leading to more efficient models. A key future direction is to extend qlty to independently partition large tensors within these scale spaces. By enabling the processing of image data across various scales, and combining the resulting outputs only at the end, qlty can help create leaner and more effective neural networks for large-scale imaging tasks.
Tracking Positional Embeddings: Aligning large images across modalities poses significant challenges, and the use of positional information is crucial in such tasks. Currently, tracking positional embeddings across subtensors can be achieved by creating a dummy tensor with positional data. However, a more efficient approach would be to enhance qlty by attaching origin and slicing parameters to each chunk. These parameters could then be used to maintain positional context and enable tensor-index to location parameters in a physical model throughout the processing pipeline, improving the accuracy and effectiveness of alignment operations.
@@ 6 Acknowledgements
We gratefully acknowledge the support of this work by the Laboratory Directed Research and Development Program of Lawrence Berkeley National Laboratory under US Department of Energy contract No. DE-AC02-05CH11231. Further support originated from the Center for Advanced Mathematics in Energy Research Applications funded via the Advanced Scientific Computing Research and the Basic Energy Sciences programs, which are supported by the Office of Science of the US Department of Energy (DOE) under contract No. DE-AC02-05CH11231, and from the National Institute of General Medical Sciences of the National Institutes of Health (NIH) under award 5R21GM129649-02. We acknowledge the use of resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility operated under Contract No. DE-AC02-05CH11231, under NERSC allocation m4055. During the preparation of this work the author used chatGPT in order to correct grammar and improve sentence construction. After using this tool/service, the author reviewed and edited the content as needed and takes full responsibility for the content of the publication.